OpenStudy (anonymous):

can someone please help my question is what is the max value for y> or equal to 3x-5 and x< or equal to -1/2x+2 and x> or equal to-5 and then it gives me f(x,y)=-3x+4y? my second question is how do you solve this system using matricies 8x+7y=5 and 4x-9y=65?

OpenStudy (anonymous):

the choices for the first question is a.-10 b.-2 c.78 d.95 for the second a.(-5,5) b.(5,65) c)(7,-7) d(65,5)

OpenStudy (anonymous):

yea sorry i went to the restroom

OpenStudy (anonymous):

\[y \ge 3x-5\] \[y \le -1/2x+2\] \[x \ge -5\]

OpenStudy (anonymous):

and then it gives me f(x,y)=-3x+4y

OpenStudy (anonymous):

oh the one on the comment sorry

OpenStudy (anonymous):

y not the x so y<orequal to -1/2x+2

OpenStudy (anonymous):

yea but idk where to go to find one whats a good website

OpenStudy (anonymous):

messed up sorry

OpenStudy (anonymous):

is it better

OpenStudy (anonymous):

nevermind i found the answer thank you for the help bye

OpenStudy (ranga):

For what it is worth, here are the steps to a similar problem: http://www.algebra.com/algebra/homework/Graphs.faq.question.548265.html